A restaurant food storage air conditioning working on air refrigeration cycle of 10 tonnes capacity works between 4 bar and 16 bar. The air leaves the cold chamber at 0oC and discharges air at 25oC to the expansion cylinder after air cooler. The expansion and compression cylinders are double acting. The mechanical efficiency of compressor and expander are 85% and 80% respectively. The compressor speed is 250 r.p.m and has a stroke of 250mm Determine: 1. C.O.P 2. Power required 3. Bore of compression and expansion cylinder (Assume isentropic compression and expansion as polytropic with n = 1.25)
1
Expert's answer
2020-04-28T12:26:33-0400

Here, It is given

"T_1=273K, T_3= 298 K, P_2=4 bar, P_1= 16 bar" , we know that one tonne of


refrigeration= "\frac{336\times1000}{24}=14000"


"\frac{T_1}{T_2}= (\frac{P_1}{P_2})^{\frac{n-1}{n}}"


"T_2= \frac{T_1}{1.32}= 207 K"


cop="\frac{T_2}{T_1-T_2}= 6.08"


W= "\frac{R}{cop}= \frac{140000}{6.08}=23026 kJ/h= 6.39 kW"